The core diameter of multimode step index fibre is 60 micrometer. The difference in refractive index is 0.013. The core refractive index is 1.46. Predict the number of guided modes when the operating wavelength is 0.75 micrometer


1
Expert's answer
2022-03-09T12:24:08-0500

"n_1=1.46\\n_2=1.46+0.013=1.473"

D=60"\mu m"

"\lambda=0.75\mu m"

NA=

"\sqrt{n_2^2-n_1^2}=\sqrt{1.473^2-1.46^2}=0.19"

NA=0.19

Number of mode

"=\frac{1}{2}\times(\frac{\pi D\times(NA)}{\lambda})^2"

Number of mode =

"=\frac{1}{2}\times(\frac{3.14\times60\times10^{-6}\times(0.19)}{0.75\times10^{-6}})^2=1200"

Number of mode=1200